OpenStudy (anonymous):

The Everton College store paid $1710 for an order of 45 calculators. The store paid $9 for each scientific calculator. The others, all graphing calculators, cost the store $55 each. How many of each type of calculator was ordered? The store ordered-scientific calculators and -graphing calculators.

OpenStudy (anonymous):

x+y=45 9x+55y=1710

OpenStudy (anonymous):

let x = $9 let y= $ 55 x+y=1710, y=1710-x 9x+55y=1710 substitution and solve